For the function f(x)=x^3 - 6x^2 - 63x +124, Determine the following:
i. The turning points
ii. Sketch the graph of the function.
iii. The x-values where the function is increasing and decreasing
1
Expert's answer
2019-09-11T09:37:40-0400

ANSWER: i. x=-3 and x=7 are the turning points.

ii. The function graph is placed in the attached file.

iii. The function decreases in the interval (-3,7). The function increases in the intervals "\quad \quad \quad \quad \left( -\infty ,-3 \right) \quad" and "\quad \left( 7,+\infty \right) \quad"

EXPLANATION: "f(x)={ x }^{ 3 }{ -6x }^{ 2 }-63x+124" , "f'(x)=3{ x }^{ 2 }{ -12x-63=3(x-7)(x+3) }" .

In the intervals "(-\infty ,-3)\quad ,(7,+\infty )\quad f'(x)>0" . Therefore, the function "y=f(x)" increases. In the interval "(-3,7)\quad \quad \quad f'(x)<0" , therefore, the function "y=f(x)" decreases.

The points x=-3 and x=7 are the turning points because at these points the derivative changes the sign.